Solve the problem. You must write two equations with two variables and show steps to solve.
Anne and Nancy use a metal alloy that is 17% copper to make jewelery. How many ounces of an alloy that is 11% copper must be mixed with an alloy that is 22% copper to form 55 ounces of the desired alloy?

Answer by jorel1380(3719)

Let n be the amount of 22% copper to be used, and m be the amount of 11% copper to be used. Then:
n+m=55 and
.22n+.11m=.17(55)
So:
n+m=55
m=55-n
.22n+.11(55-n)=9.35
.11n+6.05=9.35
.11n=3.3
n=30
m=25
They need 30 oz. of 22% alloy, and 25 oz. of 11% alloy to make their jewelery. ☺☺☺☺